Write an expression for the area of a rectangle whose length is 5 inches longer than it's width
Answer by nerdybill(7384)

Write an expression for the area of a rectangle whose length is 5 inches longer than it's width
.
Let w = width of rectangle
then from "length is 5 inches longer than it's width":
w+5 = length of rectangle
.
Since area of any rectangle is width*length:
area = w(w+5)
area = w^2+5w